OpenStudy (anonymous):

there is no graph for this. Select the coordinates of two points on the line y = -2 A. (2, -2) and (-2, 2) B. (-2,-2) and (-2, 0) C. (2, -2) and (0, -2) D. (2, -2) and (-2, -2)

OpenStudy (anonymous):

So it doesn't matter what x-coordinates are chosen, the y-coordinate will always be -2. So any point with y-coordinate of -2 will be on that line.

OpenStudy (anonymous):

yes. this question did not have a graph so.......

OpenStudy (anonymous):

You don't need one. Any point with a y-coordinate of -2 is one the line y=-2. Choose an answer where the y-coordinates are -2.

OpenStudy (anonymous):

it's either b or d?

OpenStudy (anonymous):

b has one point with y-coordinate -2 and another point with y-coordinate 0. So it can't be the correct answer.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what about d?

OpenStudy (anonymous):

Are both y-coordinates -2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great. So those two points must be on the line y=-2. But check the given answers very carefully. The way it is typed above, there are two correct answers. Are you sure you typed it correctly?

OpenStudy (anonymous):

nope. for d its ( -2, 2). i switched the numbers

OpenStudy (anonymous):

OK. Look for the answer that has y-coordinates of -2. That's the correct one.

OpenStudy (anonymous):

C? or D? i cant choose :(

OpenStudy (anonymous):

Check D very carefully. You just said you made a mistake typing it in. Are both y-coordinates -2?

OpenStudy (anonymous):

yes? these kinds of questions stress me out

OpenStudy (anonymous):

No need to stress. Answer D, as you typed it, is (2, -2) and (-2, -2). Then you said you switched the numbers and D should be (-2, 2) and (-2, -2). Am I correct?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sorry. made a mistake myself. If answer D is (-2, 2) and (-2, -2) can it be the solution?

OpenStudy (anonymous):

i am gonna say yes

OpenStudy (anonymous):

Look carefully. Are both y-coordinates -2?

OpenStudy (anonymous):

by the way you replied i am going to say no

OpenStudy (anonymous):

OK. What is the y-coordinate of (-2, 2)?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Yes. That's correct. Well done.

OpenStudy (anonymous):

yay!!!!!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

write the equation that is perpendicular to the line y = 3/2x + 4 and passes through point (-6, 3) A. y = 3/2x - 1 B. y = 3/2x + 3 C. y = -2/3x - 1 D. y = -2/3x + 3 Can you help?

OpenStudy (anonymous):

^ @ospreytriple can you help?

OpenStudy (anonymous):

Well, one characteristic of perpendicular lines is that their slopes are negative reciprocals of each other. What is the slope of the given line?

OpenStudy (anonymous):

i looked at the question and it didn't have a slope

OpenStudy (anonymous):

Sure it does, we just have to find it. The equation of the given line is y=3/2 x + 4. Recall that the general form of a linear function is y=mx+b, where m is the slope and b is the y-intercept. So what is the slope of y = 3/2 x + 4?

OpenStudy (anonymous):

ummmmmmmmmmmmmmmmmmmmmmm...... i don't know. Sorry :(

OpenStudy (anonymous):

The slope (m) is the coefficient of x, i.e. the number that is multiplying x. What number is multiplying x in y = 3/2 x + 4 ?

OpenStudy (anonymous):

1? i am not good in math.

OpenStudy (anonymous):

That's OK. What number is directly in front of the x ?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

I guess it's the way that it's typed. Let my retype it for you: \[y=\frac{ 3 }{ 2 }x\] Now, what number is multiplying the x? Hint; it's a fraction.

OpenStudy (anonymous):

the fraction that is already there? math is not my best subject. and i sound stupid

OpenStudy (anonymous):

Yes, it's the fraction that's already there.

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

So, in \[y=\frac{ 3 }{ 2 }x+4\] the x is being multiplied by \[\frac{ 3 }{ 2 }\]. You OK so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Good. 3/2 is called the coefficient of x, and it is the slope of the graph of this function. Now a line that is perpendicular to this will have a slope that is the negative reciprocal to 3/2. Do you know what that means?

OpenStudy (anonymous):

no. can you please explain it?

OpenStudy (anonymous):

No problem, it's quite simple. To take the reciprocal of a number means to turn it upside down. For example, the reciprocal of 4/5 is 5/4 and the reciprocal of 3 is 1/3. Get it? In our question we want the reciprocal of 3/2. What would that be?

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

Right. Now the word negative means change the number's sign - if it;s positive, make it negative, if it's negative, make it positive. So the negative reciprocal of 2/5 is -5/2 and the negative reciprocal of -5/6 is 6/5. We want the negative reciprocal of 3/2. What is it?

OpenStudy (anonymous):

-2/3

OpenStudy (anonymous):

Excellent. So now we know that a line that is perpendicular to \[y=\frac{ 3 }{ 2 }x+4\]must have a slope of -2/3. So, it's equation must look like\[y=-\frac{ 2 }{ 3 }+b\]. OK with that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sorry.\[y=-\frac{ 2 }{ 3 }x+b\]

OpenStudy (anonymous):

OK. To find b, we substitute the values of x and y from the given point, i.e. (-6, 3) and solve for b. Can you do that?

OpenStudy (anonymous):

sorta can you explain and i'll solve?

OpenStudy (anonymous):

OK. you'll have\[3=\left( -\frac{ 2 }{ 3 } \right)\left( -6 \right)+b\]Can you take it from here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great. What do you get for b?

OpenStudy (anonymous):

3 if that's wrong then -1

OpenStudy (anonymous):

I love that you provided two answers. The correct one is -1. So now you've got the equation as\[y=-\frac{ 2 }{ 3 }x-1\] and you're done.